OpenStudy (anonymous):

Please help me solve this equation: write in exponential form sqrt(5)xy^3z^4 with solution tnx! t.y. :D HELP!!

OpenStudy (unklerhaukus):

\[\sqrt{a}=a^{1/2}\]

OpenStudy (anonymous):

\[\sqrt{a } =a^{1/2}\] use this rule ..it will help you

OpenStudy (anonymous):

\[\sqrt{5xy ^{3}z ^{4}}=(5xy^3z^4)^{1/2}\] @ynapot97 : hope this will help you